User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair/Presenting: Jonny/Ross

In Attendance: Jonny, Ross, Tiege, Jamie, Arwin, Jim, Jesse, Bryn

Apologies: Lucas, Liz, Nick, Ellen, Lucia




  1. Meeting roster
  2. 12 hr shifts
    • Update availability page on the wiki as to if you want 12 hour shifts or not.
  3. disk deficit not on hb checklist
    • Fixed
  4. Has the slogit been edited to make summary files read only.
    • Fixed by Jamie
  5. The new script has had an update. Automatic missed scan reporting for ho 26m has now been added. You can list scans missed by commenting “t1 for start and “t2 for end, it will list all the scans missed between these two times. It Probably knows your name and the names of other observers who have done the checklist for the current experiment.
  • It currently has two bugs. Firstly, it can lose the one scan's worth of data recorded for every module change during the experiment. Secondly, it fails when the field system time is completely wrong, revert to in this case.
  1. Remember if you're on 9-5 to check if the correct KE module is loaded for the next experiment

Handover notes issues:

  1. OHIG106 (Hb): Issue with drives getting stuck - resetting HMI didn't work but eventually they came back online. The clock on HMI was out but 'reboot central' fixed the issue
  2. R1778 (KE): weather wasn't being logged (JL investigating)
    • Fixed at this stage. Caused due to power outage and a failed UPS
  3. wind measurements for Hb and Ke are stuck in the system monitor
    • Telescopes still wind stow correctly
    • Jamie to look into
/home/www/auscope/opswiki/data/pages/operations/agenda2017_02_16.txt · Last modified: 2017/02/16 02:18 by Jonathan Rogers